A newspaper article estimated that the mean cost for preventive dental care was $165 per year.
They indicated that the estimate was based on a sample of 100 people and that the margin of
error was 2.8. Assuming a 95% confidence level was used, calculate the value of the standard
error.
1
Expert's answer
2021-10-01T15:54:02-0400

Let "M" be the margin of error and "SE" be the standard error, then m is given by,

"M=Z_{\alpha/2}*SE" , Where "SE=\sigma/\sqrt{n}"

Given "M=2.8" and "Z_{\alpha/2}=Z_{0.05/2}=Z_{0.025}=1.96"

Therefore, "SE=M/Z_{0.025}=2.8/1.96=1.4286(4\displaystyle\space decimal\displaystyle\space places)."

Standard error is 1.4286.